Normal No Spac HeadingHeading 2 Font Paragraph Styles 2-97. A batch of 300 samples of rejuvenated mitochondria contains eight that are mutated (or defective). Two are selected, at random, without replacement from the batch. (a) What is the probability that the second one selected is defective given that the first one was defective? (b) What is the probability that both are defective? (c) Wltat is the probability that both are acceptable?
Answer #1

(a)

Given, the first one was defective, number of defective in remaining sample of 299 is 7.

Probability that the second one is defective = 7/299 = 0.02341137

(b)

Probability that the first one is defective = 8/300

Now, if the first one is defective, number of remaining defective samples is 7 out of remaining 299 samples

Probability that the second one is defective = 7/299

Probability that both are defective = Probability that the first one is defective * Probability that the second one is defective

= (8 / 300) * (7 / 299)

= 0.0006243032

(c)

Initial number of acceptable samples = 300 -8 = 292

Probability that the first one is acceptable = 292/300

Now, if the first one is acceptable, number of remaining acceptable samples is 291 out of remaining 299 samples

Probability that the second one is acceptable = 291 / 299

Probability that both are defective = Probability that the first one is defective * Probability that the second one is defective

= ( 292/300) * (291 / 299)

= 0.947291